# November, 2025

### v1

@@
Hinako Iseki1aaaiseki.hinako(at)st.kitasato-u.ac.jp , Shin Sasaki1bbbshin-s(at)kitasato-u.ac.jp and Kenta Shiozawa2cccshiozawa.kenta(at)kitasato-u.ac.jp 1Department of Physics, Kitasato University
@@
We study quantum mechanical wavefunctions near highly curved spaces, i.e., black holes. By utilizing the formalism developed by DeWitt, we derive the Schrödinger equations in the vicinity of the Schwarzschild and the Reissner–Nordström black hole geometries. The quantum electron cloud for the “black hydrogen atom”—an electron trapped by black holes—is particularly studied. We solve the equations and find that black holes generally attract the wavefunctions, localizing them near the horizon where the electrons are most likely to be trapped. These results imply that not only classical objects but also the quantum material and even the chemical properties of the atoms are affected by strong gravity. We also discuss black hydrogen molecules composed of multi-centered Majumdar–Papapetrou black holes.
@@ 1 Introduction
Since gravity is the weakest fundamental interaction in nature, it is usually ignored in quantum theory of elementary particles. However, it is obvious that in strong gravitational systems, such as black holes, effects of gravity can not be ignored. A typical example is the Hawking radiation of a black hole arising from the pair production/annihilation of particles in curved spacetimes [2, 3]. There are many studies on quantum natures of particles in black hole backgrounds, for example see references [4, 5, 6, 7, 8]. Very recently, quantum effects in curved spaces have been experimentally examined [9].
@@
On the other hand, the fate of the quantum matters of intermediate scales such as $1\,\text{\AA }=10^{-10}\,\mathrm{m}$ near black holes has been the subject of relatively little investigations. There are several works along this line. These include one electron atom in general curved spacetimes [10, 11], effects of gravity on hydrogen spectrum [12], a bound state of relativistic scalar particles and a rotating black hole [13], a nonrelativistic limit of scalar and Dirac fields in curved spacetimes [14, 15], gravitational effects by topological defects on the hydrogen atom [16], general relativistic effects on hydrogen systems [17]. It is no doubt that the quantum behavior of matters of atomic scale is best studied in the nonrelativistic quantum mechanics. Their quantum mechanical properties, including the detail of the material, the atomic and the chemical properties, are completely understood by the solutions to the Schrödinger equation. It is now tantalizing to study how the strong gravity effects on the quantum wavefunctions in the atomic scale. It is apparent that effects of strong gravity near a black hole can not be captured by simply introducing a Newtonian potential in the Schrödinger equation. We therefore need the Schrödinger equation that incorporates the curvature effects of spacetime.
@@
To this end, we utilize the formalism developed by DeWitt [18]. In this formalism, based on the general procedure of the analytical mechanics in the curved configuration space, the quantum mechanical Schrödinger equation in curved spaces is derived. Although this formalism subsequently developed into the quantization of gravity, including the Wheeler–DeWitt equation [19], we use it here to investigate the effects of gravity on quantum matters of atomic scale. We note that this nonrelativistic treatment is guaranteed by the local Lorentz description of the slowly moving condition. We assume that the local energy and the velocity for particles of mass $m$ obey $E_{\rm loc}\ll mc^{2}$ and $v_{\rm loc}\ll c$. We also assume that the tidal condition $|R_{\ \nu\rho\sigma}^{\mu}|\,\lambda^{2}\ll 1$ is fulfilled in the parameter ranges we consider, so that the de Broglie wavelength $\lambda$ is much shorter than the curvature radius. This remains true except within an extremely close region of the horizon. Quantum matters in the vicinity of black holes in other formalism have been studied such as quantum mechanics of constrained particles [20], quantum mechanical particles near black holes using Fermi coordinate [21, 23, 22] and so on.
@@
The organization of this paper is as follows. In the next section, we briefly introduce the Schrödinger equation in curved spaces that is developed by DeWitt [18]. In Section 3, we write down the Schrödinger equation in the Schwarzschild black hole background. We explicitly show the equations in the spherical coordinate that result from the canonical quantization based on the Cartesian coordinate. We decompose the equation by the variable separation and then study the radial equation by introducing a virtual electric potential. We show that the Schrödinger equation reduces to the confluent Heun’s differential equation, ensuring that it is solved analytically. We then solve the equation with appropriate boundary condition and show that the quantum electron cloud is mostly trapped just on the event horizon. We visualize the wavefunctions and find that the probability densities are strongly affected by the black hole. In Section 4, we discuss the charged black holes. We study the quantum wavefunctions in the backgrounds of the non-extremal, extremal and over-extremal Reissner–Nordström black holes. The wavefunctions in the regions inside the inner (Cauchy) and outside the outer (event) horizons are studied. We find that the famous gravitational repulsion force near the core of the Reissner–Nordström black hole is observed in the wavefunction. The extremal limit and the over-extremal cases are also discussed. In Section 5, we discuss the black hole molecule that consists of the two extremal black hole nuclei and one electron. This corresponds to the hydrogen molecular ion $\mathrm{H}_{2}^{+}$. We show that in a specific region of parameters, the equations can be solved analytically. Section 6 is devoted to conclusion and discussions. We discuss a possible realization of the black hole atoms in the present Universe.
@@ 2 Schrödinger equation in curved space
where $g_{ij}$ is a spatial metric of the configuration space spanned by the coordinates $q^{i}$ and $V(q)$ is a potential. Here $\dot{q}^{i}=\frac{{\mathrm{d}}q^{i}}{{\mathrm{d}}t}$ and $t$ is a parameter that specifies the particle motion. The vector potential $a_{i}$ should be included when we consider background magnetic fields. The Riemann and the Ricci curvature tensors associated with the spatial metric $g_{ij}$ are defined as
@@
where $p_{i}$ is the canonical momentum conjugate to $q^{i}$. The quantum Hamiltonian $\hat{H}_{\pm}$ receives the curvature and the metric corrections [^1] Note that there is a potential ambiguity on the quantum Hamiltonian. We have two equivalent choices $\hat{H}_{\pm}=\hat{H}\pm\frac{\hbar^{2}}{12}R^{(3)}$ that are consistent with the classical limit $\hbar\to 0$. We here employ $\hat{H}_{+}$ rather than $\hat{H}_{-}$ for definiteness. This kind of non-uniqueness of quantum theory in curved space is often recognized. See [24] for example. :
@@
In order to ensure the conservation of the probability, the $t$-derivative of the wavefunction $\psi$ is modified in the curved space:
@@
Here the Hamiltonian $\hat{H}_{+}$ is given in (2.4). For later convenience, we rewrite the equation in the explicit form. In the following, we consider $a_{i}=0$ and $q^{i}$ as the Cartesian coordinate $x^{i}$ ($i=1,2,3$) for a three-dimensional space. Then the momentum is given by
@@
The curvature independent part of the Hamiltonian $\hat{H}$ in (2.4) is evaluated as
@@
Note that the derivative $\hat{A}_{i}$ acts on the all parts of its right side including the wavefunction. By distributing the $\hat{A}_{i}$ to the relevant terms and collect the terms altogether, we have the explicit form of the Schrödinger equation:
@@
In the following, we focus on the stationary state for $\psi$ and assume that the background metric $g_{ij}$ is independent of $t$. In this case, we have $\frac{\partial g}{\partial t}=0$ and the wavefunction is decomposed as $\psi(t,x)=\mathrm{e}^{-i\frac{E}{\hbar}t}\psi(x)$ where $E$ is a constant. Then the equation for $\psi(x)$ is found to be
@@ 3 Quantum wavefunction near Schwarzschild black hole
where $r^{2}=x^{2}+y^{2}+z^{2}$ and $a=\frac{2G_{\mathrm{N}}M}{c^{2}}$ is a constant. A three-dimensional spatial geometry arises from the time slice of the metric (3.1). The point $r=a$ corresponds to the event horizon that the spatial metric diverges. The spatial metric in the Cartesian coordinate in the matrix notation is given by
@@
Then we have $g=\det g_{ij}=\frac{r}{r-a}$ and $R^{(3)}=0$. We stress that we should begin with the Cartesian coordinate rather than the radial one since the canonical quantization condition (2.3) is justified only in the former coordinate system. We obtain the correct quantum mechanical equation in the spherical coordinate $(r,\theta,\phi)$ by staring from (3.1) in the $(x,y,z)$ system and then switch to the $(r,\theta,\phi)$ system at the final stage of the calculations.
After tedious calculations, we find that the equation (2.11) in the $(r,\theta,\phi)$ coordinate is given by
@@
The wavefunction is decomposed as $\psi(r,\theta,\phi)=R(r)Y_{l}^{m}(\theta,\phi)$. Then we find that the equations for each part are given by
@@
where $l=0,1,2,\ldots$ are constants. The equation for $Y^{m}_{l}$ is the same with the one in the flat space. Then $Y_{l}^{m}(\theta,\phi)$ is found to be the spherical harmonics:
@@ 3.1 Free particle near black hole
where we have assumed $E>0$ and defined $k^{2}=\frac{2\mu E}{\hbar^{2}}$. We can show that the equation (3.12) has regular singular points at $r=0$ and $r=a$. The indicial equation of (3.12) reveals that the solutions at $r=a$ behaves like $R(r)\sim(r-a)^{0}$ and $(r-a)^{\frac{1}{2}}$. Then they are regular at the horizon but the derivative $R^{\prime}(r)$ diverges in general. Since the radial coordinate $r$ in (3.1) becomes timelike in $r<a$ and thus it ceases to be a suitable coordinate for the wavefunction, we look for smooth solutions outside the horizon $r>a$. The Schwarzschild solution is asymptotically flat and solutions to the equation (3.12) must approach the ones in the flat space. It is easy to show that in the asymptotic flat regime $r\to\infty$, the equation has the solution given by the spherical Bessel functions that behave as $R(r)\sim\frac{1}{r}\mathrm{e}^{\pm ikr}$. This fact allows us to employ the boundary condition $\psi\to 0$ ($r\to\infty$).
For the consistency, we examine the equation in the asymptotic limit $r\gg a$. In this region, the equation is compared to that in the flat space. After rearranging the coefficient of $R^{\prime\prime}$ to be unity and by the rescaling [^2] This rescaling arises from matching the integral measure to that of the flat space $\sqrt{g}|R|^{2}\to|R|^{2}$. Namely, $R\to g^{-\frac{1}{4}}R=\left(1-\frac{a}{r}\right)^{-\frac{1}{4}}R\sim\left(1+\frac{a}{4r}\right)R$. $\hat{R}=\left(1+\frac{a}{4r}\right)R$, the equation results in the form
@@
We now rewrite the equation (3.12). By introducing the dimensionless coordinate $x=\frac{r}{a}$ and substitute $R(x)=\mathrm{e}^{\sigma x}H(x)$, the equation (3.12) becomes the following form:
$H^{\prime\prime}(x)+\left\{\frac{\frac{3}{2}}{x}+\frac{-\frac{1}{2}}{x-1}+2\sigma\right\}H^{\prime}(x)+\frac{1}{x(x-1)}\left\{(2\sigma+a^{2}k^{2})x-\frac{3}{2}\sigma-l(l+1)\right\}H(x)=0,$ (3.14)
@@
where we have set $\sigma=\pm iak$. We find that (3.14) is the confluent form of the Heun’s differential equation. The parameters [^3] The standard form of the Heun’s differential equation of confluent type is $H^{\prime\prime}(z)+\left[\frac{\gamma}{z}+\frac{\delta}{z-1}+\varepsilon\right]H^{\prime}(z)+\frac{\alpha z-q}{z(z-1)}H(z)=0.$ are $\gamma=\frac{3}{2}$, $\delta=-\frac{1}{2}$, $\varepsilon=2\sigma=\pm 2iak$, $\alpha=2\sigma+a^{2}k^{2}$ and $q=-\frac{3}{2}\sigma-l(l+1)$. The general solution is therefore given by $R(r)=\mathrm{e}^{\pm ikr}H(r)$, where $H(r)$ is the confluent Heun function defined by the parameters given above. The signs $\pm$ correspond to the outgoing and ingoing modes, respectively. It is physically preferable to choose the ingoing mode as the solution. The confluent Heun function has two linearly independent asymptotic expansion at $x\to\infty$: $H(x)\sim x^{-\alpha/\epsilon}$ and $H(x)\sim\mathrm{e}^{-\epsilon x}x^{-\gamma-\delta+\alpha/\epsilon}$ [25]. The physically relevant solution $\psi\to 0$ corresponds to the second branch. In order to see this behavior, it is useful to visualize the wavefunction.
@@
To this end, we solve the equation numerically and make plots. The probability density $\sqrt{g}|R(r)|^{2}r^{2}$ in the radial direction is shown in Fig. 1.
@@
Figure 1: The probability density $\sqrt{g}|R(r)|^{2}r^{2}$ in the radial direction. Comparison between the flat space and the black hole background cases for $a=1,l=1$ with $k=1$ (left), $k=5$ (middle) and $k=10$ (right). The normalization factors are appropriately chosen. The red dotted line corresponds to the event horizon $r=a$.
@@
We find that the wavefunction is most strongly attracted near the horizon. At the same time, the wavefunction exhibits overall shifts toward the black hole direction. Note that although $R(r)$ itself is finite at $r=a$, the probability density diverges at the horizon due to the factor $\sqrt{g}=\sqrt{\frac{r}{r-a}}$. It is also observed that the higher the energy, the more it is attracted toward the center. This can be confirmed by comparing the distance between any two peaks in the wavefunction. As the energy increases by $k=1,5,10$, the distances between any adjacent peaks become smaller. This is natural since the energy is the source of gravity. These characteristic behaviors indicate that quantum particles are also strongly affected by gravity. The 3D plots in Fig. 2 clearly show how gravity affects the wavefunction.
@@
Figure 2: The 3D plots for the probability density $|\psi(r,\theta,\phi)|^{2}$ for free particle. The cases of flat space (left) and the black hole background (right). The parameters are fixed to $a=1$, $l=1$ and $k=1$. The event horizon is shown in the black sphere. The wavefunction is affected by the black hole.

### v2

@@
Hinako Iseki1aaaiseki.hinako(at)st.kitasato-u.ac.jp, Shin Sasaki1bbbshin-s(at)kitasato-u.ac.jp, and Kenta Shiozawa2cccshiozawa.kenta(at)kitasato-u.ac.jp 1Department of Physics, Kitasato University
@@
We study quantum mechanical wave functions near highly curved spaces, i.e., black holes. By utilizing the formalism developed by DeWitt, we derive the Schrödinger equations in the vicinity of the Schwarzschild and the Reissner–Nordström black hole geometries. The quantum electron cloud for the “black hydrogen atom”—an electron trapped by black holes—is particularly studied. We solve the equations and find that black holes generally attract the wave functions, localizing them near the horizon where the electrons are most likely to be trapped. These results imply that not only classical objects but also the quantum material and even the chemical properties of the atoms are affected by strong gravity. We also discuss black hydrogen molecules composed of multicentered Majumdar–Papapetrou black holes.
@@ 1 Introduction
Since gravity is the weakest fundamental interaction in nature, it is usually ignored in quantum theory of elementary particles. However, it is obvious that in strong gravitational systems, such as black holes, effects of gravity cannot be ignored. A typical example is the Hawking radiation of a black hole arising from the pair production/annihilation of particles in curved spacetimes [2, 3]. There are many studies on quantum natures of particles in black hole backgrounds, for example, see Refs. [4, 5, 6, 7, 8]. Very recently, quantum effects in curved spaces have been experimentally examined [9].
@@
On the other hand, the fate of the quantum matters of intermediate scales such as $1\,\text{\AA }=10^{-10}\,\mathrm{m}$ near black holes has been the subject of relatively few investigations. There are several works along this line. These include one electron atom in general curved spacetimes [10, 11], effects of gravity on the hydrogen spectrum [12], a bound state of relativistic scalar particles and a rotating black hole [13], a nonrelativistic limit of scalar and Dirac fields in curved spacetimes [14, 15], gravitational effects by topological defects on the hydrogen atom [16], and general relativistic effects on hydrogen systems [17]. There is no doubt that the quantum behavior of matters of atomic scale is best studied in the nonrelativistic quantum mechanics. Their quantum mechanical properties, including the detail of the material, the atomic, and the chemical properties, are completely understood by the solutions to the Schrödinger equation. It is now tantalizing to study how the strong gravity affects on the quantum wave functions in the atomic scale. It is apparent that effects of strong gravity near a black hole cannot be captured by simply introducing a Newtonian potential in the Schrödinger equation. The Newtonian approximation is valid when a quantum system is far from the black hole, namely, $r\gg r_{\text{H}}$, where $r_{\text{H}}$ is the radius of the black hole event horizon. In order to examine the strong gravitational effects on quantum systems near the horizon at $r\sim r_{\text{H}}$, where the Newtonian approximation is violated, the Schrödinger equation that incorporates curvature effects of spacetime is required.
@@
To this end, we utilize the formalism developed by DeWitt [18]. In this formalism, based on the general procedure of the analytical mechanics in the curved configuration space, the quantum mechanical Schrödinger equation in curved spaces is derived. Although this formalism subsequently developed into the quantization of gravity, including the Wheeler–DeWitt equation [19], we use it here to investigate the effects of gravity on quantum matters of atomic scale. We note that this nonrelativistic treatment is guaranteed by the local Lorentz description of the slowly moving condition. We assume that the local energy and the velocity for particles of mass $m$ obey $E_{\rm loc}\ll mc^{2}$ and $v_{\rm loc}\ll c$. We also assume that the tidal condition $|R_{\ \nu\rho\sigma}^{\mu}|\,\lambda^{2}\ll 1$ is fulfilled in the parameter ranges we consider, so that the de Broglie wavelength $\lambda$ is much shorter than the curvature radius. This means that the energy scale of quantum particles is smaller than that of the gravitational tidal forces. The condition allows us to utilize the nonrelativistic approximation. Throughout this paper, we focus on black holes of atomic scale. Then the tidal force condition remains true except within an extremely close region of the horizon where $|R^{\mu}{}_{\nu\rho\sigma}|^{-\frac{1}{2}}$ and $\lambda$ are of the same order of magnitude, i.e., the Bohr radius of the hydrogen. Therefore, we are interested in the region $r_{\text{H}}<r<\alpha\,r_{\text{H}}$, where $\alpha$ is an $\mathcal{O}(10^{0}\sim 10^{1})$ parameter. Quantum matters in the vicinity of black holes in other formalisms have been studied, such as quantum mechanics of constrained particles [20], quantum mechanical particles near black holes using Fermi coordinates [21, 23, 22], and so on.
@@
The organization of this paper is as follows. In the next section, we briefly introduce the Schrödinger equation in curved spaces that is developed by DeWitt [18]. In Sec. 3, we write down the Schrödinger equation in the Schwarzschild black hole background. We explicitly show the equations in the spherical coordinate that result from the canonical quantization based on the Cartesian coordinate. We decompose the equation by the variable separation and then study the radial equation by introducing a virtual electric potential. We show that the Schrödinger equation reduces to the confluent Heun’s differential equation, ensuring that it is solved analytically. We then solve the equation with an appropriate boundary condition and show that the quantum electron cloud is mostly trapped just on the event horizon. We visualize the wave functions and find that the probability densities are strongly affected by the black hole. In Sec. 4, we discuss the charged black holes. We study the quantum wave functions in the backgrounds of the nonextremal, extremal, and overextremal Reissner–Nordström black holes. The wave functions in the regions inside the inner (Cauchy) and outside the outer (event) horizons are studied. We find that the famous gravitational repulsion force near the core of the Reissner–Nordström black hole is observed in the wave function. The extremal limit and the overextremal cases are also discussed. In Sec. 5, we discuss the black hole molecule that consists of the two extremal black hole nuclei and one electron. This corresponds to the hydrogen molecular ion $\mathrm{H}_{2}^{+}$. We show that in a specific region of parameters, the equations can be solved analytically. Sec. 6 is devoted to the conclusion and discussions. We discuss a possible realization of the black hole atoms in the present Universe.
@@ 2 Schrödinger equation in curved space
where $g_{ij}$ is a spatial metric of the configuration space spanned by the coordinates $q^{i}$, and $V(q)$ is a potential. Here, $\dot{q}^{i}=\frac{{\mathrm{d}}q^{i}}{{\mathrm{d}}t}$, and $t$ is a parameter that specifies the particle motion. The vector potential $a_{i}$ should be included when we consider background magnetic fields. The Riemann and the Ricci curvature tensors associated with the spatial metric $g_{ij}$ are defined as
@@
where $p_{i}$ is the canonical momentum conjugate to $q^{i}$. The quantum Hamiltonian $\hat{H}_{\pm}$ receives the curvature and the metric corrections [^1] Note that there is a potential ambiguity on the quantum Hamiltonian. We have two equivalent choices $\hat{H}_{\pm}=\hat{H}\pm\frac{\hbar^{2}}{12}R^{(3)}$ that are consistent with the classical limit $\hbar\to 0$. We here employ $\hat{H}_{+}$ rather than $\hat{H}_{-}$ for definiteness. This kind of nonuniqueness of quantum theory in curved space is often recognized. See [24], for example. ,
@@
In order to ensure the conservation of the probability, the $t$-derivative of the wave function $\psi$ is modified in the curved space,
@@
Here, the Hamiltonian $\hat{H}_{+}$ is given in (2.4). For later convenience, we rewrite the equation in the explicit form. In the following, we consider $a_{i}=0$ and $q^{i}$ as the Cartesian coordinate $x^{i}$ ($i=1,2,3$) for a three-dimensional space. Then, the momentum is given by
@@
The curvature-independent part of the Hamiltonian $\hat{H}$ in (2.4) is evaluated as
@@
Note that the derivative $\hat{A}_{i}$ acts on all parts of its right side including the wave function. By distributing the $\hat{A}_{i}$ to the relevant terms and collecting the terms altogether, we have the explicit form of the Schrödinger equation,
@@
In the following, we focus on the stationary state for $\psi$ and assume that the background metric $g_{ij}$ is independent of $t$. In this case, we have $\frac{\partial g}{\partial t}=0$, and the wave function is decomposed as $\psi(t,x)=\mathrm{e}^{-i\frac{E}{\hbar}t}\psi(x)$, where $E$ is a constant. Then, the equation for $\psi(x)$ is found to be
@@ 3 Quantum wavefunction near Schwarzschild black hole
where $r^{2}=x^{2}+y^{2}+z^{2}$, and $a=\frac{2G_{\mathrm{N}}M}{c^{2}}$ is a constant. A three-dimensional spatial geometry arises from the time slice of the metric (3.1). The point $r=a$ corresponds to the event horizon that the spatial metric diverges. The spatial metric in the Cartesian coordinate in the matrix notation is given by
@@
Then we have $g=\det g_{ij}=\frac{r}{r-a}$ and $R^{(3)}=0$. We stress that we should begin with the Cartesian coordinate rather than the radial one since the canonical quantization condition (2.3) is justified only in the former coordinate system. We obtain the correct quantum mechanical equation in the spherical coordinate $(r,\theta,\phi)$ by starting from (3.1) in the $(x,y,z)$ system and then switch to the $(r,\theta,\phi)$ system at the final stage of the calculations.
After tedious calculations, we find that Eq. (2.11) in the $(r,\theta,\phi)$ coordinate is given by
@@
The wave function is decomposed as $\psi(r,\theta,\phi)=R(r)Y_{l}^{m}(\theta,\phi)$. Then, we find that the equations for each part are given by
@@
where $l=0,1,2,\ldots$ are constants. The equation for $Y^{m}_{l}$ is the same as the one in the flat space. Then, $Y_{l}^{m}(\theta,\phi)$ is found to be the spherical harmonics
@@ 3.1 Free particle near black hole
where we have assumed $E>0$ and defined $k^{2}=\frac{2\mu E}{\hbar^{2}}$. We can show that Eq. (3.12) has regular singular points at $r=0$ and $r=a$. The indicial equation of (3.12) reveals that the solutions at $r=a$ behave like $R(r)\sim(r-a)^{0}$ and $(r-a)^{\frac{1}{2}}$. Then, $R(r)$ is regular but $R^{\prime}(r)$ for $R(r)\sim(r-a)^{\frac{1}{2}}$ diverges at the horizon. Since the radial coordinate $r$ in (3.1) becomes timelike in $r<a$ and thus it ceases to be a suitable coordinate for the wave function, we look for smooth solutions outside the horizon $r>a$. The Schwarzschild solution is asymptotically flat, and solutions to Eq. (3.12) must approach the ones in the flat space. It is easy to show that in the asymptotic flat regime $r\to\infty$, the equation has the solution given by the spherical Bessel functions that behave as $R(r)\sim\frac{1}{r}\mathrm{e}^{\pm ikr}$.
For consistency, we examine the equation in the asymptotic limit $r\gg a$. In this region, the equation is compared to that in the flat space. After rearranging the coefficient of $R^{\prime\prime}$ to be unity and by the rescaling [^2] This rescaling arises from matching the integral measure to that of the flat space $\sqrt{g}|R|^{2}\to|R|^{2}$. Namely, $R\to g^{-\frac{1}{4}}R=\left(1-\frac{a}{r}\right)^{-\frac{1}{4}}R\sim\left(1+\frac{a}{4r}\right)R$. $\hat{R}=\left(1+\frac{a}{4r}\right)R$, the equation results in the form
@@
We now rewrite Eq. (3.12). By introducing the dimensionless coordinate $x=\frac{r}{a}$ and substituting $R(x)=\mathrm{e}^{\sigma x}H(x)$, Eq. (3.12) becomes the following form:
$H^{\prime\prime}(x)+\left\{\frac{\frac{3}{2}}{x}+\frac{\frac{1}{2}}{x-1}+2\sigma\right\}H^{\prime}(x)+\frac{1}{x(x-1)}\left\{(2\sigma+a^{2}k^{2})x-\frac{3}{2}\sigma-l(l+1)\right\}H(x)=0,$ (3.14)
@@
where we have set $\sigma=\pm iak$. We find that (3.14) is the confluent form of Heun’s differential equation. The parameters [^3] The standard form of Heun’s differential equation of confluent type is $H^{\prime\prime}(z)+\left[\frac{\gamma}{z}+\frac{\delta}{z-1}+\varepsilon\right]H^{\prime}(z)+\frac{\alpha z-q}{z(z-1)}H(z)=0.$ are $\gamma=\frac{3}{2}$, $\delta=\frac{1}{2}$, $\varepsilon=2\sigma=\pm 2iak$, $\alpha=2\sigma+a^{2}k^{2}$, and $q=-\frac{3}{2}\sigma-l(l+1)$. The general solution is, therefore, given by $R(r)=\mathrm{e}^{\pm ikr}H(r)$, where $H(r)$ is the confluent Heun function defined by the parameters given above. The signs $\pm$ correspond to the outgoing and ingoing modes, respectively. The confluent Heun function has two linearly independent asymptotic expansions at $x\to\infty$: $H(x)\sim x^{-\alpha/\varepsilon}$ and $H(x)\sim\mathrm{e}^{-\varepsilon x}x^{-\gamma-\delta+\alpha/\varepsilon}$ [25]. The physically relevant solution that behaves as $R(r)\sim\frac{1}{r}\mathrm{e}^{\pm ikr}$ ($r\to\infty$) corresponds to the first branch. Physically, this implies that a purely outgoing or ingoing mode at infinity originates from the scattering of waves at the horizon. The local analysis at $r=a$ gives $R\sim A+B\sqrt{r-a}$, where $A,B$ are constants. A boundary condition at infinity, which we will take such that $R\sim\frac{\sin kr}{r}$ is a real-valued function, fixes a particular linear combination of these two local behaviors through the global properties of the confluent Heun function, or equivalently, through a numerical integration.
@@
In order to see the behavior of the wave function near the black hole, it is useful to visualize the solution. To this end, we solve the equation numerically and make plots. The probability density $\sqrt{g}|R(r)|^{2}r^{2}$ in the radial direction is shown in Fig. 1.
(a) Radial probability density of the $l=1$.
(b) Radial probability density of the $l=2$.
@@
Figure 1: The probability density $\sqrt{g}|R(r)|^{2}r^{2}$ in the radial direction. Comparison between the flat space and the black hole background cases for $a=1$. (a) $l=1$ with $k=1$ (left), $k=5$ (middle), and $k=10$ (right), and (b) $l=2$ with $k=1$ (left), $k=5$ (middle), and $k=10$ (right). The red dotted line corresponds to the event horizon $r=a$. The infinity is set to be $r_{\text{max}}=1000$ and we define the boundary condition $R(r_{\text{max}})=0$ to find the numerical solutions with Mathematica. The normalization factors are appropriately chosen.
@@
We find that the wave function is most strongly attracted near the horizon. At the same time, the wave function exhibits overall shifts toward the black hole direction. Note that although $R(r)$ itself is finite at $r=a$, the probability density diverges at the horizon due to the factor $\sqrt{g}=\sqrt{\frac{r}{r-a}}$. It is also observed that the higher the energy, the more it is attracted toward the center. This can be confirmed by comparing the distance between any two peaks in the wave function. As the energy increases by $k=1,5,10$, the distances between any adjacent peaks become smaller. A naive interpretation of this fact is that objects with higher energy (mass) are attracted more strongly by gravity. These characteristic behaviors indicate that quantum particles are also strongly affected by gravity. The 3D plots in Fig. 2 clearly show how gravity affects the wave function. Note that we employ real-valued solutions for $R(r)$. This means that the linear combination of the ingoing and outgoing modes produces the standing wave, resulting in the vanishing of the radial current of probability $j_{r}\propto R\partial_{r}\bar{R}-\bar{R}\partial_{r}R=0$.
@@
Figure 2: The 3D plots for the probability density $|\psi(r,\theta,\phi)|^{2}$ for free particles. The cases of flat space (left) and the black hole background (right). The parameters are fixed to $a=1$, $l=1$, and $k=1$. The values increase from blue to red and the event horizon is shown in the black sphere.